Alberto is 6 years older than Luis. Within 4 years the sum of their ages will be 58 years old calculate they have now
Answer by Boreal(15235)

Luis's current age is x, Alberto's x+6
in 4 years, x+4+x+10=58
2x+14=58
2x=44
x=22 Luis ANSWER
x+6=28 Alberto ANSWER
in 4 years, 26 and 32, which is 58.